THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 98 — 1 April 2009

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues, It is my pleasure to present you the 98^{th} issue of the Magellanic Clouds Newsletter. There is lots to entertain you, from galactic structure and initial mass functions, star clusters, variable stars and X-ray sources, to dust from supernovae and evolved stars — as well as a re-examination of my favourite star (besides the Sun), WOH G64. The next issue will be distributed on the 7^{th} of June 2009; the deadline for contributions is the 6^{th} of June.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

The Optical Gravitational Lensing Experiment. OGLE-III Photometric Maps of the Small Magellanic Cloud

A. Udalski^{1}, I. Soszyński^{1}, M.K. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, Ł. Wyrzykowski^{3}, O. Szewczyk^{2}, K. Ulaczyk^{1} and R. Poleski^{1}

1Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2Universidad de Concepción, Departamento de Fisica, Casilla 160–C, Concepción, Chile 3Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK

We present OGLE-III Photometric Maps of the Small Magellanic Cloud. They contain precise, calibrated VI photometry of about 6.2 million stars from 41 OGLE-III fields in the SMC observed regularly in the years 2001–2008 and covering about 14 square degrees in the sky. Also precise astrometry of these objects is provided. One of the fields, SMC140, is centered on the 47 Tucanae Galactic globular cluster providing unique data on this object. We discuss quality of the data and present a few color–magnitude diagrams of the observed fields. All photometric data are available to the astronomical community from the OGLE Internet archive.

Published in Acta Astronomica 58, 329 (2008) Available from arXiv:0901.4632 and from http://acta.astrouw.edu.pl/Vol58/n4/pap58 4 3.pdf

Structure of the SMC — Stellar component distribution from 2MASS data

I. Gonidakis ^{1}, E. Livanou^{1}, E. Kontizas^{2}, U. Klein^{3}, M. Kontizas^{1}, M. Belcheva^{1}, P. Tsalmantza^{4} and A. Karampelas^{1}

1Department of Astrophysics Astronomy & Mechanics, Faculty of Physics, University of Athens, GR-15783 Athens, Greece 2Institute for Astronomy and Astrophysics, National Observatory of Athens, P.O. Box 20048, GR-11810 Athens, Greece 3Radioastronomisches Institut der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 4Max-Planck-Institut für Astronomie, Konigstuhl 17, 69117 Heidelberg, Germany

The spatial distribution of the SMC stellar component is investigated from 2MASS data. The morphology of the different age populations is presented. The center of the distribution is calculated and compared with previous estimations. The rotation of the stellar content and possible consequence of the presence of dark matter is discussed. The different stellar populations are identified through a CMD diagram of the 2MASS data. Isopleth contour maps are produced in each case, to reveal the spatial distribution. The derived density profiles are discussed. The older stellar population follows an exponential profile at projected diameters of about 5 kpc (∼ 5^{∘}) for the major axis and ∼ 4 kpc for the minor axis, centred at RA: 0^{h}51^{m}, Dec: −73^{∘}7′ (J2000.0). The centre coordinates are found to be the same for all the different age population maps and are in good accordance with the kinematical centre of the SMC. However they are found to be considerably different to the coordinates of the centre of the gas distribution. The fact that the older population is found in an exponential disk suggests that the stellar content is rotating, a possible consequence of dark matter presence. The strong interactions between the MCs and the MilkyWay might explain the difference in the distributions of the stellar and gas components. The lack of an observed velocity element, which implies an absence of rotation and contradicts the consequences of an exponential profile of the stellar component, may also be a result of gravitational interactions.

Accepted for publication in A&A Available from arXiv:0812.0880

Spectroscopic Signatures of the Superorbital Period in the Neutron Star Binary LMC X-4

Joseph Neilsen^{1,2}, Julia C. Lee^{1,2}, Michael A. Nowak^{3}, Konrad Dennerl^{4} and Saeqa Dil Vrtilek^{2}

1Harvard University, Department of Astronomy, USA 2Harvard-Smithsonian Center for Astrophysics, USA 3MIT-Kavli Center for Astrophysics and Space Research, USA 4Max-Planck-Institut für Extraterrestrische Physik, Germany

We present the first high-resolution X-ray study of emission line variability with superorbital phase in the neutron star binary LMC X-4. Our analysis provides new evidence from X-ray spectroscopy confirming accretion disk precession as the origin of the superorbital period. The spectra, obtained with the Chandra High-Energy Transmission Grating Spectrometer (HETGS) and the XMM-Newton Reflection Grating Spectrometer (RGS), contain a number of emission features, including lines from hydrogen-like and helium-like species of N, O, Ne, and Fe, a narrow O vii RRC, and fluorescent emission from cold Fe. We use the narrow RRC and the Heα triplets to constrain the temperature and density of the (photoionized) gas. By comparing spectra from different superorbital phases, we attempt to isolate the contributions to line emission from the accretion disk and the stellar wind. There is also evidence for highly ionized iron redshifted and blueshifted by ∼ 25, 000 km s^{−1}. We argue that this emission originates in the inner accretion disk, and show that the emission line properties in LMC X-4 are natural consequences of accretion disk precession.

Accepted for publication in ApJ Available from arXiv:0902.0786

The complete Initial Mass Function down to the subsolar regime in the Large Magellanic Cloud with Hubble Space Telescope ACS Observations

Nicola Da Rio^{1}, Dimitrios Gouliermis^{1} and Thomas Henning^{1}

1Max-Planck-Institut für Astronomie, Heidelberg, Germany

In this photometric study of the stellar association LH 95 in the Large Magellanic Cloud (LMC) we focus on the premain Sequence (PMS) population in order to construct, for the first time, the sub-solar initial mass function (IMF) in the LMC. We use the deepest photometry ever performed in the LMC with the Advanced Camera for Surveys (ACS) on-board the Hubble Space Telescope (HST). We carry out a Monte Carlo technique to subtract the contribution of the general field of LMC and we isolate the central region in the observed area of the association. We study the mass function of its field-subtracted population. For this purpose, we introduce a new set of evolutionary models, derived from the calculations on the evolution of PMS stars by Siess et al. We use these models with our observations of LH 95 to derive the IMF of the system, which is reliably constructed down to 0.43 M_{⊙}, the lowest mass ever observed within reasonable completeness in the Magellanic Clouds. Consequently, its construction offers an outstanding improvement in our understanding of the low-mass star formation in the LMC. The system IMF of LH 95 shows a definite change in its slope at 1 M_{⊙}, where it becomes more shallow. In general, the shape of this IMF agrees very well with the average Galactic IMF, down to the sub-solar regime. As far as the slope of this system IMF is concerned, it is found to be somewhat more shallow than the corresponding classical Galactic IMF in the sub-solar regime, probably due to unresolved binarity, while for stars with M > 1 M_{⊙} it becomes slightly steeper. We do not find significant differences in the shape of the overall IMF of LH 95 from that of each of the three individual sub-clusters, suggesting that the IMF of LH 95 is not subject to local variability.

Accepted for publication in ApJ Available from arXiv:0902.0758

Upper Limits on Pulsed Radio Emission from the 6.85 s X-ray Pulsar XTE J0103−728 in the Small Magellanic Cloud

Fronefield Crawford^{1}, Duncan R. Lorimer^{2}, Brian M. Devour^{1}, Brian P. Takacs^{1} and Vladislav I. Kondratiev^{2,3,4}

1Department of Physics and Astronomy, Franklin and Marshall College, P.O. Box 3003, Lancaster, PA 17604, USA 2Department of Physics, West Virginia University, Morgantown, WV 26506, USA 3National Radio Astronomy Observatory, P.O. Box 2, Green Bank, WV 24944, USA 4Astro Space Center of the Lebedev Physical Institute, Profsoyuznaya str. 84/32, Moscow 117997, Russia

X-ray pulsations with a 6.85 s period were recently detected in the Small Magellanic Cloud (SMC) and were subsequently identified as originating from the Be/X-ray binary system XTE J0103−728. The recent localization of the source of the X-ray emission has made a targeted search for radio pulsations from this source possible. The detection of pulsed radio emission from XTE J0103−728 would make it only the second system after PSR B1259−63 that is both a Be/X-ray binary and a radio pulsar. We observed XTE J0103−728 in Feb 2008 with the Parkes 64-m radio telescope soon after the identification of the source of X-ray pulsations was reported in order to search for corresponding radio pulsations. We used a continuous 6.4 hour observation with a 256 MHz bandwidth centered at 1390 MHz using the center beam of the Parkes multibeam receiver. In the subsequent data analysis, which included a folding search, a Fourier search, a fast-folding algorithm search, and a single-pulse search, no pulsed signals were found for trial dispersion measures (DMs) between 0 and 800 pc cm^{−3}. This DM range easily encompasses the expected values for sources in the SMC. We place an upper limit of ∼ 45 mJy kpc^{2} on the luminosity of periodic radio emission from XTE J0103−728 at the epoch of our observation, and we compare this limit to a range of luminosities measured for PSR B1259−63, the only Be/X-ray binary currently known to emit radio pulses. We also compare our limit to the radio luminosities of neutron stars having similarly long spin periods to XTE J0103−728. Since the radio pulses from PSR B1259−63 are eclipsed and undetectable during the portion of the orbit near periastron, repeated additional radio search observations of XTE J0103−728 may be valuable if it is undergoing similar eclipsing and if such observations are able to sample the orbital phase of this system well.

Accepted for publication in Astrophysical Journal Available from arXiv:0902.0995

Could the compact remnant of SN 1987A be a quark star?

T.C. Chan^{1}, K.S. Cheng^{1}, T. Harko^{1}, H.K. Lau^{2}, L.M. Lin^{2}, W.M. Suen^{3} and X.L. Tian^{1}

1Department of Physics and Center for Theoretical and Computational Physics, The University of Hong Kong, Hong Kong, P.R. China 2Department of Physics and Institute of Theoretical Physics, The Chinese University of Hong Kong, Hong Kong, P.R. China 3McDonnell Center for the Space Sciences, Department of Physics, Washington University, St. Louis, USA

The standard model for Type II supernovae explosion, confirmed by the detection of the neutrinos emitted during the supernova explosion, predicts the formation of a compact object, usually assumed to be a neutron star. However, the lack of the detection of a neutron star or pulsar formed in the SN 1987A still remains an unsolved mystery. In this paper we suggest that the newly formed neutron star at the center of SN 1987A may undergo a phase transition after the neutrino trapping time scale (∼ 10 s). Consequently the compact remnant of SN 1987A may be a strange quark star, which has a softer equation of state than that of neutron star matter. Such a phase transition can induce the stellar collapse and result in a large amplitude stellar oscillations. We use a three dimensional Newtonian hydrodynamic code to study the time evolution of the temperature and density at the neutrinosphere. Extremely intense pulsating neutrino fluxes, with submillisecond period and with neutrino energy (> 30 MeV) can be emitted because the oscillations of the temperature and density are out of phase almost 180^{∘}. If this is true we predict that the current X-ray emission from the compact remnant of SN 1987A will be lower than 10^{34} erg s^{−1}, and it should be a thermal bremsstrahlung spectrum for a bare strange star with surface temperature of around ∼ 10^{7} K.

Accepted for publication in ApJ Available from arXiv:0902.0653

Star complexes and stellar populations in NGC 6822 — Comparison with the Magellanic Clouds

A. Karampelas^{1}, A. Dapergolas^{2}, E. Kontizas^{2}, E. Livanou^{1}, M. Kontizas^{1}, I. Bellas-Velidis^{2} and J. M. Vílchez^{3}

1Department of Astrophysics, Astronomy & Mechanics, Faculty of Physics, University of Athens, Greece 2Institute for Astronomy and Astrophysics, National Observatory of Athens, Greece 3Instituto de Astrofísica de Andalucía (CSIC), Spain

The star complexes (large scale star forming regions) of NGC 6822 were traced and mapped and their size distribution was compared with the size distribution of star complexes in the Magellanic Clouds (MCs). Furthermore, the spatial distributions of different age stellar populations were compared with each other. The star complexes of NGC 6822 were determined by using the isopleths, based on star counts, of the young stars of the galaxy, using a statistical cutoff limit in density. In order to map them and determine their geometrical properties, an ellipse was fitted to every distinct region satisfying this minimum limit. The Kolmogorov-Smirnov statistical test was used to study possible patterns in their size distribution. Isopleths were also used to study the stellar populations of NGC 6822. The star complexes of NGC 6822 were detected and a list of their positions and sizes was produced. Indications of hierarchical star formation, in terms of spatial distribution, time evolution and preferable sizes were found in NGC 6822 and the MCs. The spatial distribution of the various age stellar populations has indicated traces of an interaction in NGC 6822, dated before 350 ± 50 Myr.

Accepted for publication in A&A Available from arXiv:0902.1934

The Magellanic Bridge as a DLA System: Physical Properties of Cold Gas toward PKS 0312−770

Toru Misawa^{1,2}, Jane C. Charlton^{1}, Henry A. Kobulnicky^{3}, Bart P. Wakker^{4} and Joss Bland-Hawthorn^{5}

1Department of Astronomy & Astrophysics, The Pennsylvania State University, University Park, PA 16802, USA 2Cosmic Radiation Laboratory, RIKEN, 2-1 Hirosawa, Wako, Saitama 351-0198, Japan 3Department of Physics & Astronomy, University of Wyoming, Laramie, WY 82070, USA 4Department of Astronomy, University of Wisconsin-Madison, 475 North Charter Street, Madison, WI 53706, USA 5School of Physics, University of Sydney, NSW 2006, Australia

We measure the physical properties of a local multi-component absorption-line system at V_{⊙} ∼ 200 km s^{−1} toward the quasar PKS 0312−077 behind the Magellanic Bridge (MB) using Hubble Space Telescope STIS spectroscopy in conjunction with photoionization modeling. At an impact parameter of ∼ 10 kpc from the Small Magellanic Cloud (SMC), this sightline provides a unique opportunity to probe the chemical properties and ionization structure in a nearby absorption line system with a column density of log N_{H I} ∼ 20.2, at the transition between Damped Lyman Alpha (DLA) and sub-DLA systems. We find that metallicity of −1.0 < log(Z/Z_{⊙}) < −0.5 and ionization parameter of −6 < log U < −5 for three low-ionization components and log U ∼ −2.6 for one high-ionization component. One component at V_{⊙} = 207 km s^{−1} shows an α-element abundance log(Si/H) ∼ −5.0, making it ∼ 0.2 dex more metal rich than both SMC H ii regions and stars within the MB and the SMC. The N/Si ratio in this component is log(N/Si) = −0.3 ± 0.1, making it comparable to other N-poor dwarf galaxies and ∼ 0.2 dex lower than H ii regions in the SMC. Another component at V_{⊙} = 236 km s^{−1} shows a similar Si/H ratio but has log(N/Si) = −1.0 ± 0.2, indicating a nitrogen deficiency comparable to that seen in the most N-poor DLA systems. These differences imply different chemical enrichment histories between components along the same sightline. Our results suggest that, if these absorbers are representative some fraction of DLA systems, then 1) DLA systems along single sight-lines do not necessarily represent the global properties of the absorbing cloud, and 2) the chemical composition within a given DLA cloud may be inhomogeneous.

Accepted for publication in ApJ Available from arXiv:0902.0208

Molecular and Atomic Gas in the Large Magellanic Cloud — I. Conditions for CO Detection

T. Wong^{1,2}, A. Hughes^{3,2}, Y. Fukui^{4}, A. Kawamura^{4}, N. Mizuno^{4,5}, J. Ott^{6,7}, E. Muller^{4,2}, J.L. Pineda^{8}, D.E. Welty^{1}, S. Kim^{9}, Y. Mizuno^{4}, M. Murai^{4} and T. Onishi^{4}

1University of Illinois, USA 2CSIRO ATNF, Australia 3Swinburne University, Australia 4Nagoya University, Japan 5ALMA-J Project, NAOJ, Japan 6NRAO, USA 7Caltech Astronomy, USA 8JPL, Caltech, USA 9Sejong University, South Korea

We analyze the conditions for detection of CO(1–0) emission in the Large Magellanic Cloud (LMC), using the recently completed second NANTEN CO survey. In particular, we investigate correlations between CO integrated intensity and H i integrated intensity, peak brightness temperature, and line width at a resolution of 2.6′ (∼ 40 pc). We find that significant H i column density (exceeding ∼ 10^{21} cm^{−2}) and peak brightness temperature (exceeding ∼ 20 K) are necessary but not sufficient conditions for CO detection, with many regions of strong H i emission not associated with molecular clouds. The large scatter in CO intensities for a given H i intensity persists even when averaging on scales of > 200 pc, indicating that the scatter is not solely due to local conversion of H i into H_{2} near GMCs. We focus on two possibilities to account for this scatter: either there exist spatial variations in the I(CO) to N(H_{2}) conversion factor, or a significant fraction of the atomic gas is not involved in molecular cloud formation. A weak tendency for CO emission to be suppressed for large H i linewidths supports the second hypothesis, insofar as large linewidths may be indicative of warm H i, and calls into question the likelihood of forming molecular clouds from colliding H i flows. We also find that the ratio of molecular to atomic gas shows no significant correlation (or anti-correlation) with the stellar surface density, though a correlation with midplane hydrostatic pressure P_{h} is found when the data are binned in P_{h}. The latter correlation largely reflects the increasing likelihood of CO detection at high H i column density.

Accepted for publication in ApJ Available from arXiv:0902.1578

The stellar ancestry of supernova progenitors in the Magellanic Clouds — I. The most recent supernovae in the Large Magellanic Cloud

Carles Badenes^{1}, Jason Harris^{2}, Dennis Zaritsky^{3} and José Luis Prieto^{4}

1Princeton University, USA 2NOAO, USA 3University of Arizona, USA 4Ohio State University, USA

We use the star formation history map of the Large Magellanic Cloud recently published by Harris & Zaritsky to study the sites of the eight smallest (and presumably youngest) supernova remnants in the Cloud: SN 1987A, N 158A, N 49, and N 63A (core collapse remnants), 0509−67.5, 0519−69.0, N 103B, and DEM L71 (Type Ia remnants). The local star formation histories provide unique insights into the nature of the supernova progenitors, which we compare with the properties of the supernova explosions derived from the remnants themselves and from supernova light echoes. We find that corcollapse supernovae are always associated with vigorous star formation in the recent past. In the case of SN 1987A, the time of the last peak of star formation (12 Myr) matches very well the lifetime of a star with the mass of its blue supergiant progenitor (20 M_{⊙}). More recent peaks of star formation can lead to supernovae with more massive progenitors, which opens the possibility of a Type Ib/c origin for SNRs N 158A and N 63A. Stars more massive than 21.5 M_{⊙} are very scarce around SNR N 49, implying that the magnetar SGR 0526−66 in this SNR was either formed elsewhere or came from a progenitor with a mass well below the 30 M_{⊙} threshold suggested in the literature. Three of our four Ia SNRs are associated with old, metal poor stellar populations. This includes SNR 0509−67.5, which is known to have been originated by an extremely bright Type Ia event, and yet is located very far away from any sites of recent star formation, in a population with a mean age of 7.9 Gyr. On the other hand, the Type Ia SNR N 103B is indeed associated with recent star formation, and might have had a relatively younger and more massive progenitor with substantial mass loss before the explosion. We discuss these results in the context of our present understanding of core collapse and Type Ia supernova progenitors.

Submitted to ApJ Available from arXiv:0902.2787

Photometric analysis of Magellanic Cloud R Coronae Borealis Stars

Robyn M. Woollands^{1}, P.L. Cottrell^{1} and A. Udalski^{2}

1Dept of Physics & Astronomy, University of Canterbury, Christchurch, New Zealand 2Warsaw University Observatory, Poland

This paper presents the initial results of a multi-site photometric programme to examine the extraordinary behaviour displayed by 18 R Coronae Borealis (RCB) stars in the Magellanic Clouds (MCs). RCB stars exhibit a unique variability whereby they undergo rapid declines of up to several magnitudes. These are thought to be caused by the formation of dust in the stellar environment which reduces the brightness. The monitoring programme comprised the collection of UBV RI photometric data using five telescopes located at three different southern hemisphere longitudes (Las Campanas Observatory in Chile, Mount Joun University Observatory in New Zealand, and the Southern African Large Telescope (SALT) in South Africa). Examination of the data acquired in the V and I filters resulted in the identification of a total of 18 RCB declines occurring in four stars. Construction of colour–magnitude diagrams (V vs. V–I), during the recovery to maximum light were undertaken in order to study the unique colour behaviour associated with the RCB declines. The combined recovery slope for the four stars was determined to be 3.37 ± 0.24, which is similar to the value of 3.1 ± 0.1 calculated for galactic RCB stars (Skuljan et al. 2003). These results may imply that the nature of the dust (i.e. the particle size) is similar in both our Galaxy and the MCs.

Accepted for publication in Publications of the Astronomical Society of Australia Available from arXiv:0902.1212

Looking outside the Galaxy: the discovery of chemical anomalies in 3 old Large Magellanic Cloud clusters

A. Mucciarelli^{1}, L. Origlia^{2}, F.R. Ferraro^{1} and E. Pancino^{2}

1Dipartimento di Astronomia, Universitá di Bologna, Italy 2INAF - Osservatorio Astronomico di Bologna, Italy

By using the multifiber spectrograph FLAMES mounted at the ESO-VLT, we have obtained high-resolution spectra for 18 giant stars, belonging to 3 old globular clusters of the Large Magellanic Cloud (namely NGC 1786, 2210 and 2257). While stars in each cluster showed quite homogeneous iron content, within a few cents of dex (the mean values being Fe/H]= −1.75±0.01 dex, −1.65±0.02 dex and −1.95±0.02 dex for NGC 1786, 2210 and 2257, respectively), we have detected significant inhomogeneities for the [Na/Fe], [Al/Fe], [O/Fe] and [Mg/Fe] abundance ratios, with evidence of [O/Fe] vs. [Na/Fe] and [Mg/Fe] vs. [Al/Fe] anticorrelations. The trends detected nicely agree with those observed in Galactic Globular Clusters, suggesting that such abundance anomalies are ubiquitous features of old stellar systems and they do not depend on the parent galaxy environment. In NGC 1786 we also detected two extreme O-poor, Na-rich stars. This is the first time that a firm signature of extreme chemical abundance anomalies has been found in an extragalactic stellar cluster.

Accepted for publication in ApJ Letters Available from arXiv:0902.4778

Measuring Dust Production in the Small Magellanic Cloud Core-Collapse Supernova Remnant 1E 0102.2−7219

Karin M. Sandstrom^{1}, Alberto D. Bolatto^{2}, Snežana Stanimirović^{3}, Jacco Th. van Loon^{4} and J.D.T. Smith^{5}

1Astronomy Department, University of California, Berkeley, USA 2Department of Astronomy and Laboratory for Millimeter-wave Astronomy, University of Maryland, College Park, USA 3Department of Astronomy, University of Wisconsin, Madison, USA 4Astrophysics Group, Lennard-Jones Laboratories, Keele University, UK 5Ritter Astrophysical Research Center, University of Toledo, USA

We present mid-infrared spectral mapping observations of the core-collapse supernova remnant 1E 0102.2−7219 in the Small Magellanic Cloud using the InfraRed Spectrograph (IRS) on the Spitzer Space Telescope. The remnant shows emission from fine structure transitions of neon and oxygen as well as continuum emission from dust. Comparison of the mid-IR dust emission with observations at x-ray, radio and optical wavelengths shows that the dust is associated with the supernova ejecta and is thus newly formed in the remnant. The spectrum of the newly formed dust is well reproduced by a model that includes 3 × 10^{−3} M_{⊙} of amorphous carbon dust at 70 K and 2 × 10^{−5} M_{⊙} of Mg_{2}SiO_{4} (forsterite) at 145 K. Our observations place a lower limit on the amount of dust in the remnant since we are not sensitive to the cold dust in the unshocked ejecta. We compare our results to observations of other core-collapse supernovae and remnants, particularly Cas A where very similar spectral mapping observations have been carried out. We observe a factor of ∼ 10 less dust in E 0102 than seen in Cas A, although the amounts of amorphous carbon and forsterite are comparable. Finally, we present evidence suggesting that the grain size distribution of the newly formed dust in E 0102 has been altered by the hot plasma behind the reverse shock.

Accepted for publication in Astrophysical Journal Available from arXiv:0810.2803

The global gas and dust budget of the Large Magellanic Cloud: AGB stars and supernovae, and the impact on the ISM evolution

M. Matsuura ^{1,2}, M.J. Barlow^{2}, A.A. Zijlstra^{3}, P.A. Whitelock^{4,5}, M.-R.L. Cioni^{6}, M.A.T. Groenewegen^{7}, K. Volk^{8}, F. Kemper ^{3}, T. Kodama^{1}, E. Lagadec^{3}, M. Meixner^{9}, G.C. Sloan^{10} and S. Srinivasan^{11}

1National Astronomical Observatory of Japan, Japan 2University College London, UK 3University of Manchester, UK 4South African Astronomical Observatory, South Africa 5University of Cape Town, South Africa 6University of Hertfordshire, UK 7Royal Observatory of Belgium, Belgium 8Gemini Observatory, USA 9Space Telescope Science Institute, USA 10Cornell University, USA 11Johns Hopkins University, USA

We report on an analysis of the gas and dust budget in the the interstellar medium (ISM) of the Large Magellanic Cloud (LMC). Recent observations from the Spitzer Space Telescope enable us to study the mid-infrared dust excess of asymptotic giant branch (AGB) stars in the LMC. This is the first time we can quantitatively assess the gas and dust input from AGB stars over a complete galaxy, fully based on observations. The integrated mass-loss rate over all intermediate and high mass-loss rate carbon-rich AGB candidates in the LMC is 8.5 × 10^{−3} M_{⊙} yr^{−1}, up to 2.1 × 10^{−2} M_{⊙} yr^{−1}. This number could be increased up to 2.7 × 10^{−2} M_{⊙} yr^{−1}, if oxygen-rich stars are included. This is overall consistent with theoretical expectations, considering the star formation rate when these low- and intermediate-mass stars where formed, and the initial mass functions. AGB stars are one of the most important gas sources in the LMC, with supernovae (SNe), which produces about 2–4×10^{−2} M_{⊙} yr^{−1}. At the moment, the star formation rate exceeds the gas feedback from AGB stars and SNe in the LMC, and the current star formation depends on gas already present in the ISM. This suggests that as the gas in the ISM is exhausted, the star formation rate will eventually decline in the LMC, unless gas is supplied externally. Our estimates suggest ‘a missing dust-mass problem’ in the LMC, which is similarly found in high-z galaxies: the accumulated dust mass from AGB stars and possibly SNe over the dust life time (400–800 Myrs) is significant less than the dust mass in the ISM. Another dust source is required, possibly related to star-forming regions.

Accepted for publication in MNRAS Available from arXiv:0903.1123

The Mass-Loss Return From Evolved Stars to the LMC: Empirical Relations for Excess Emission at 8 and 24 μm

Sundar Srinivasan^{1}, Margaret Meixner^{2}, Claus Leitherer^{2}, Uma Vijh^{2}, Kevin Volk^{3}, Robert D. Blum^{4}, Brian L. Babler^{5}, Miwa Block^{6}, Steve Bracker^{5}, Martin Cohen^{7}, Charles W. Engelbracht^{6}, Bi-Qing For^{6}, Karl D. Gordon^{2},

Jason Harris^{4}, Joseph L. Hora^{8}, Remy Indebetouw^{9}, Francisca Markwick-Kemper^{10}, Marilyn Meade^{5}, Karl A. Misselt^{6}, Marta Sewilo^{2} and Barbara Whitney^{11}

1Johns Hopkins University, USA 2Space Telescope Science Institute, USA 3Gemini Observatory, USA 4National Optical Astronomy Observatory, USA 5University of Wisconsin at Madison, USA 6University of Arizona, USA 7University of California at Berkeley, USA 8Harvard-Smithsonian Center for Astrophysics, USA 9University of Virginia, USA 10University of Manchester, UK 11Space Science Institute, USA

We present empirical relations describing excess emission from evolved stars in the Large Magellanic Cloud (LMC) using data from the Spitzer Space Telescope SAGE (Surveying the Agents of a Galaxy’s Evolution) survey which includes the IRAC 3.6, 4.5, 5.8 and 8.0 μm and MIPS 24, 70 and 160 μm bands. We combine the SAGE data with the Two Micron All Sky Survey (2MASS; J, H and K_{s}) and the optical Magellanic Cloud Photometric Survey (MCPS; U, B, V and I) point source catalogs in order to create complete spectral energy distributions (SEDs) of the asymptotic giant branch (AGB) star candidates in the LMC. AGB star outflows are among the main producers of dust in a galaxy, and this mass loss results in an excess in the fluxes observed in the 8 and 24 μm bands. The aim of this work is to investigate the mass-loss return by AGB stars to the interstellar medium of the LMC by studying the dependence of the infrared excess flux on the total luminosity. We identify oxygen-rich, carbon-rich and extreme AGB star populations in our sample based on their 2MASS and IRAC colors. The SEDs of oxygen- and carbon-rich AGB stars are compared with appropriate stellar photosphere models to obtain the excess flux in all the IRAC bands and the MIPS 24 μm band. Extreme AGB stars are dominated by circumstellar emission at 8 and 24 μm thus we approximate their excesses with the flux observed in these bands. We find about 16,000 O-rich, 6300 C-rich and 1000 extreme sources with reliable 8 μm excesses, and about 4500 O-rich, 5300 C-rich and 960 extreme sources with reliable 24 μm excesses. The excesses are in the range 0.1 mJy – 5 Jy. The 8 and 24 μm excesses for all three types of AGB candidates show a general increasing trend with luminosity. The color temperature of the circumstellar dust derived from the ratio of the 8 and 24 μm excesses decreases with an increase in excess, while the 24 μm optical depth increases with excess. The extreme AGB candidates are the major contributors to the mass loss, and we estimate the total AGB mass-loss return to the LMC to be (5.9–13)×10^{−3} M_{⊙} yr^{−1}.

Accepted for publication in Astronomical Journal Available from arXiv:0903.1661

The Physical Properties of the Red Supergiant WOH G64: The Largest Star Known?

Emily M. Levesque^{1}, Philip Massey^{2}, Bertrand Plez^{3} and Knut A.G. Olsen^{4}

1Institute for Astronomy, University of Hawaii, 2680 Woodlawn Dr., Honolulu, HI 96822, USA 2Lowell Obsevatory, 1400 West Mars Hill Road, Flagstaff, AZ 86001, USA 3GRAAL, Université Montpellier, CNRS, 34095 Montpellier, France 4National Optical Astronomy Observatory, 950 North Cherry Avenue, Tucson, AZ 85748, USA

WOH G64 is an unusual red supergiant (RSG) in the Large Magellanic Cloud (LMC), with a number of properties that set it apart from the rest of the LMC RSG population, including a thick circumstellar dust torus, an unusually late spectral type, maser activity, and nebular emission lines. Its reported physical properties are also extreme, including the largest radius for any star known and an effective temperature that is much cooler than other RSGs in the LMC, both of which are at variance with stellar evolutionary theory. We fit moderate-resolution optical spectrophotometry of WOH G64 with the MARCS stellar atmosphere models, determining an effective temperature of 3400 ± 25 K. We obtain a similar result from the star’s broadband V–K colors. With this effective temperature, and taking into account the flux contribution from the aysmmetric circumstellar dust envelope, we calculate log(L/L_{⊙}) = 5.45 ± 0.05 for WOH G64, quite similar to the luminosity reported by Ohnaka and collaborators based on their radiative transfer modeling of the star’s dust torus. We determine a radius of R/R_{⊙} = 1540, bringing the size of WOH G64 and its position on the H-R diagram into agreement with the largest known Galactic RSGs, although it is still extreme for the LMC. In addition, we use the Ca ii triplet absorption feature to determine a radial velocity of 294 ± 2 km s^{−1} for the star; this is the same radial velocity as the rotating gas in the LMC’s disk, which confirms its membership in the LMC and precludes it from being an unusual Galactic halo giant. Finally, we describe the star’s unusual nebula emission spectrum; the gas is nitrogen-rich and shock-heated, and displays a radial velocity that is significantly more positive than the star itself by 50 km s^{−1}.

Accepted for publication in The Astronomical Journal Available from arXiv:0903.2260

Population Parameters of Intermediate-Age Star Clusters in the Large Magellanic Cloud. I. NGC 1846 and its Wide Main Sequence Turnoff

Paul Goudfrooij^{1}, Thomas H. Puzia^{2}, Vera Kozhurina-Platais^{1} and Rupali Chandar^{3}

1STScI, USA 2HIA, Canada 3University of Toledo, USA

The Advanced Camera for Surveys on board the Hubble Space Telescope has been used to obtain deep, high-resolution images of the intermediate-age star cluster NGC 1846 in the Large Magellanic Cloud. We present new color–magnitude diagrams (CMDs) based on F435W, F555W, and F814W imaging. We test the previously observed broad main sequence turnoff region for ”contamination” by field stars and (evolved) binary star systems. We find that while these impact the number of objects in this region, none can fully account for the large color spread. Our results therefore solidify the recent finding that stars in the main sequence turnoff region of this cluster have a large spread in color which is unrelated to measurement errors or contamination by field stars, and likely due to a ∼ 300 Myr range in the ages of cluster stars. An unbiased estimate of the stellar density distribution across the main sequence turnoff region shows that the spread is fairly continuous rather than strongly bimodal as suggested previously. We fit the CMDs with several different sets of theoretical isochrones, and determine systematic uncertainties for population parameters when derived using any one set of isochrones. We note a degeneracy between age and [α/Fe], which can be lifted by matching the shape (curvature) of the full red giant branch in the CMD. We find that stars in the upper part of the main sequence turnoff region are more centrally concentrated than those in any other region of the CMD, including more massive red giant branch and asymptotic giant branch stars. We consider several possible formation scenarios which account for the unusual features observed in the CMD of NGC 1846.

Accepted for publication in The Astronomical Journal Available from arXiv:0903.4511

Double-Overtone Cepheids in the Large Magellanic Cloud

W.A. Dziembowski^{1,2} and R. Smolec^{2}

1Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warsaw, Poland 2Copernicus Astronomical Center, ul. Bartycka 18, 00-716 Warsaw, Poland

One of the most interesting results from the OGLE-III study of the LMC Cepheids is the large number of objects that pulsate simultaneously in the first and second overtone (denoted 1O/2O). Double-mode Cepheids yield important constraint on stellar evolution models. We show that great majority of the LMC 1O/2O Cepheids have masses M = 3.0 ± 0.5 M_{⊙}. According to current stellar evolution calculations, these masses are lower than needed for the blue loop in the helium burning phase to reach the instability strip. On the other hand, we found most of these stars significantly overluminous if they are crossing the instability strip before helium ignition. A possible solution of this discrepancy is to allow for a large overshooting from the convective core in the main sequence phase. We also discuss origin of double-mode pulsation. At the short period range we find two types of resonances that are conducive to this form of pulsation. However, at longer periods, it has a different (non-resonant) origin.

Accepted for publication in Acta Astronomica Available from arXiv:0903.2676

The Optical Gravitational Lensing Experiment. The OGLE-III Catalog of Variable Stars. III. RR Lyrae Stars in the Large Magellanic Cloud

I. Soszyński^{1}, A. Udalski^{1}, M.K. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, Ł. Wyrzykowski^{3}, O. Szewczyk^{2}, K. Ulaczyk^{1} and R. Poleski^{1}

1Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2Universidad de Concepción, Departamento de Fisica, Casilla 160–C, Concepción, Chile 3Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK

The third part of the OGLE-III Catalog of Variable Stars comprises 24 906 RR Lyr stars in the Large Magellanic Cloud (LMC). This sample consists of 17 693 fundamental-mode (RRab), 4958 first-overtone (RRc), 986 double-mode (RRd) and 1269 suspected second-overtone (RRe) pulsators. 66 objects are foreground Galactic RR Lyr stars. The catalog data include basic photometric and astrometric properties of the RR Lyr stars, multi-epoch VI photometry and finding charts. We detected one new RR Lyr star with additional eclipsing variations. The spatial distribution of RR Lyr stars in the LMC is distinctly non-spherical and it is elongated in the same direction as the LMC bar. The basic statistical features of RR Lyr stars in the LMC are provided. The apparent V-band magnitudes for RRab stars have the modal value at 19.36 mag, and for overtone RR Lyr stars it is about 19.32 mag. The mean periods for RRab, RRc and RRe stars are 0.576, 0.337 and 0.270 days, respectively.

Accepted for publication in Acta Astronomica Available from arXiv:0903.2482

The initial mass function of the rich young cluster NGC 1818 in the Large Magellanic Cloud

Q. Liu^{1,2}, R. de Grijs^{3,1}, L.C. Deng^{1}, Y. Hu^{1,2}, I. Baraffe^{4} and S.F. Beaulieu^{5}

1National Astronomical Observatories, Chinese Academy of Sciences, Beijing 100012, P.R. China 2Graduate University of the Chinese Academy of Sciences, Beijing 100049, P.R. China 3Department of Physics & Astronomy, The University of Sheffield, Sheffield S3 7RH, UK 4CRAL, École Normale Supérieure, 46 allée d’Italie, 69007 Lyon, France 5Département de Physique, de Génie Physique et d’Optique and Centre de Recherche en Astrophysique du Québec, Université Laval, Québec, QC G1V 0A6, Canada

We use deep Hubble Space Telescope photometry of the rich, young (∼20–45 Myr-old) star cluster NGC 1818 in the Large Magellanic Cloud to derive its stellar mass function (MF) down to ∼ 0.15 M_{⊙}. This represents the deepest robust MF thus far obtained for a stellar system in an extragalactic, low-metallicity ([Fe/H] ≃ −0.4 dex) environment. Combining our results with the published MF for masses above 1.0 M_{⊙}, we obtain a complete present-day MF. This is a good representation of the cluster’s initial MF (IMF), particularly at low masses, because our observations are centred on the cluster’s uncrowded half-mass radius. Therefore, stellar and dynamical evolution of the cluster will not have affected the low-mass stars significantly. The NGC 1818 IMF is well described by both a lognormal and a broken power-law distribution with slopes of Γ = 0.46 ± 0.10 and Γ ≃ −1.35 (Salpeter-like) for masses in the range from 0.15 to 0.8 M_{⊙} and greater than 0.8 M_{⊙}, respectively. Within the uncertainties, the NGC 1818 IMF is fully consistent with both the Kroupa solar-neighbourhood and the Chabrier lognormal mass distributions.

Accepted for publication in MNRAS Available from arXiv:0903.4787

Conference Papers

The Physical Properties of Red Supergiants

Emily M. Levesque^{1}

1Institute for Astronomy, University of Hawaii, 2680 Woodlawn Dr., Honolulu, HI 96822, USA

Red supergiants (RSGs) are a He-burning phase in the evolution of moderately massive stars (10–25 M_{⊙}). For many years, the assumed physical properties of these stars placed them at odds with the predictions of evolutionary theory. We have recently determined new effective temperatures and luminosities for the RSG populations of galaxies with a factor of ∼ 8 range in metallicity, including the Milky Way, the Magellanic Clouds, and M 31. We find that these new physical properties greatly improve the agreement between the RSGs and the evolutionary tracks, although there are still notable difficulties with modeling the physical properties of RSGs at low metallicity. We have also examined several unusual RSGs, including VY CMa in the Milky Way, WOH G64 in the LMC, and a sample of four RSGs in the Magellanic Clouds that show considerable variations in their physical parameters, most notably their effective temperatures. For all of these stars we reexamine their placement on the H-R diagram, where they have appeared to occupy the ”forbidden” region to the right of the Hayashi track. We have updated current understanding of the physical properties of VY CMa and WOH G64; in the case of the unusual Magellanic Cloud variables, we conclude that these stars are undergoing an unstable evolutionary phase not previously associated with RSGs.

Oral contribution, published in Hot and Cool: Bridging Gaps in Massive Star Evolution Available from arXiv:0902.2789

More On Pulsating B-type stars in the Magellanic Clouds

P.D. Diago^{1}, J. Gutiérrez-Soto^{2,3}, J. Fabregat^{1} and C. Martayan^{2,4}

1Observatori Astronòmic de la Universitat de València. Edif. Instituts d’Investigació. Pol. La Coma, 46980 Paterna, València, Spain 2GEPI, Observatoire de Paris, CNRS, Université Paris Diderot, Place Jules Janssen 92195 Meudon Cedex, France 3LESIA, Observatoire de Paris, CNRS, Université Paris Diderot, Place Jules Janssen 92195 Meudon Cedex, France 4Royal Observatory of Belgium, 3 Avenue Circulaire, B-1180 Brussels, Belgium

We present here the results of our research for B-type pulsators in low metallicity environments, searching for shortterm periodic variability in a large sample of B and Be stars in the Magellanic Clouds (MC), for which the fundamental astrophysical parameters were accurately determined. A significant number of β Cephei and SPB-like pulsators at lowmetallicity have been detected, conflicting with the current theoretical models of pulsation. In addition, we have placed these pulsating stars in the HR diagram mapping the observational instability regions for the MC metallicities. The large sample of B and Be stars analysed allows us to make a reliable statistics of the pulsating B-type stars in the MC. Finally, we have made a comparison between pulsational theory and observations in low metallicity environments.

Oral contribution, published in the 38^{th} LIAC/HELAS-ESTA/BAG, Liège (Belgium) July 2008 Available from arXiv:0902.4354

High excitation nebulae around Magellanic Wolf-Rayet stars

Manfred W. Pakull^{1}

1Observatoire Astronomíque de Strasbourg, France

The SMC harbours a class of hot nitrogen-sequence Wolf-Rayet stars (WNE) that display only relatively weak broad emission lines. This indicates low mass-loss rates and makes them also hard to detect. However, such stars are possible emitters of strong He^{+} Lyman continua that in turn could ionize observable He ii regions, i.e. highly excited H ii regions emitting nebular He ii 4686 emission. We here report the discovery of a rare HeIII region in the SMC wich is located OB association NGC 249 around the weak-lined WN star SMC WR 10. WR 10 is of particularly interesting since it is a single star showing the presence of atmospheric hydrogen. While analysing the spectrum in the framework of two popular WR atmosphere models, we found for the same input parameters strongly discrepant predictions (by 1 dex) of the He^{+} Lyman continuum. A second aspect of the work reported here concerns the beautiful MCELS images which clearly reveal a class of strongly [O iii] 5007 emitting (blue-coded) nebulae. Not unexpectedly, most of the ’blue’ nebulae are known Wolf-Rayet bubbles, but new bubbles around a few WRs are also detected. Moreover, we report the existence of blue nebulae without associated known WRs and discuss the possibility that they reveal weak-wind WR stars with very faint stellar He ii 4686 emission. Alternatively, such nebulae might hint at the hitherto missing population of relatively low-mass, hot He stars predicted by massive binary evolution calculations. Such a binary system is probably responsible for the ionization of the unique He ii 4686-emitting nebula N 44C.

Oral contribution, published in IAU Symposium 256; ”The Magellanic System: Stars, Gas, and Galaxies”, Jacco Th. van Loon & Joana M. Oliveira (eds.) Available from arXiv:0903.3563